Problem 1: Isoquant, Isocost Cost Minimizing Approach to Factor Selection: Suppose that as part of the UNCCCC Paris Agreemente that the houthe nate di ve ponthie slopes of any tangenthi ve effional labohnical subse d) Plot and label this isoquant onTP, and MP, curves and provide a brief explanation. i) Is the MP, increasing, constant, diminishing, or decreasing returns toTables and Figures: Table 1: Table 2: workers/hr) Crobots/hr) TC ($/hr) L-intercept K-intercept workers/h (robots/hr) 10 60 2Table 3: TPL(L, KSR) MPL(L, KSR) TPL(L, KLR) MPL(L, KSR) (worker (e-buses/hr) (e-buses/ (e-buses/hr) (e-buses/ s/hr) worker h
Answer #1

.. we conside the production fraction Q(K, L) = AL25 k3/5 a i human capital workys training , entunship Programmis etc. . i२१ - ks (२१)% ___० (*) Sk ___२५3 -) . : . iso quart. CS Scanned with main on Tu as the equation ofL= 5, : 93.9 | k: 293 : = 52.8 . 243. 243 = When L= 10 ; k=- | Jun L=20, 2 -3 , (26/3 k = 22:243 - 32.8 24 20) =20, k= up = 4Table 1 workns/h 1 ans 148) K (robots/h corobores/ho] 83.8 52.8 10 20 32.8 40 60 20.8 15.9 13.1 .80 11.3 100 120 1.40 10 16014 ,k 83 . 8 به .امه أ۱۶۰۰ . 13.1 ج 3Sر دو 11 1 مه ° { 0 s 10 20 10 50 0 0 10 20 140 150 L |cs Scanned with CamScanneres es Marginal substitution rate of technical Manila etsa umpk 315k215 MPK . . Hne Q (k; 2) = A2 25 K3/5 Hue A=5 Q (KL = SL 32 now homly rate of labor W=&al hom homly opportunity cost of hining a robot. 03 $3 ) non : Now stor I socost equa Tez WLtokK-in Ancept of iso cost equation Tsubstitute L=0] : - 120 = 0.+ 3k .=> K= 40 than by joining and ik in Ancept, we get the isowha Te=300 L-inducept of iso cost line [ k=0] 300 = L + (340) => 22360 K-in incept of iso cost line [ 2=0] 300 = 0+ 3k -> K=2Go - 20. 150 + 1 AO IN Q3s 20 5 10| 150 6010 120 150 180 200 210 250 2309 | 20 | . . 이 등 by The noe of hiring labori 쀼 응 r 를